5. Consider the demand and supply functions for the notebooks market.

𝑄𝐷=10,000βˆ’100𝑝 𝑄𝑆=900𝑝


a. Make a table with the corresponding supply and demand schedule.

b. Draw the corresponding graph.

c. Is it possible to find the price and quantity of equilibrium with the graph method?

d. Find the price and quantity of equilibrium by solving the system of equations.


1
Expert's answer
2022-03-23T14:02:59-0400

A)

Given the Qd=10,000-100p and Qs=900p


Price Quantity demanded Quantity supplied

0 10,000 0

5 9,500 4,500

10 9,000 9,000

15 8,500 13,500

20 8,000 18,000

B) Graph



C) From the graph Equilibrium (P,Q) IS

P=$10

Q=9,000


D)

At equilibrium Qd=Qs ,where Qd=10,000-100P and Qs=900P

10000-100P=900P

1000P=10,000

P=$10


Qd=10,000-100p

=10,000-(100"\times"10)

=10,000-1000

Q=9,000

Therefore at equilibrium P=$10,Q=900